OpenStudy (anonymous):

The side lengths of a square (L) are increasing at a constant rate of 2 cm/min. If the area of a square is A. Then dA/dL is equal to?

OpenStudy (anonymous):

A= L^2 ===> dA/dL= 2L *dL/dt === > 2L *2 = 4L

OpenStudy (anonymous):

the answer is 2L

OpenStudy (anonymous):

oh yeah its right sorry lol i didn't see your s da/dl

OpenStudy (anonymous):

ok , great

OpenStudy (anonymous):

it depends on the law of area

OpenStudy (anonymous):

so the area of a square is L x B so thats why its L^2

OpenStudy (anonymous):

since B = L in the square, it's better to write the area in this form A=L*L =L^2

OpenStudy (anonymous):

i see thank you very much :)

OpenStudy (anonymous):

yoy are welcome